Problem 17-47 Derive Amounts for Profit Variance Analysis (LO 17-5) Classics, Ltd., details cars. Classics wants to compare this quarters results with those for last quarter, which is believed to be typical for operations. Assume that the following information is provided: Last Quarter This Quarter Number of detailings Variable costs Contribution margin 560 763 $75,040 29,120 S45,920 S71,200 35,520 535,680 Required: Compute the flexible budget and sales activity variance and prepare a profit variance analysis. (Hint: Use last quarter as the master budget and this quarter as actual.) Do not round intermediate calculations. Indicate the effect of each variance by selecting F for favorable, or U for unfavorable. If there is no effect, do not select either option.) Variable Cost Variance Flexible Bud Sales Activity Variance Master Bud Actual Sales Price Variance Sales revenue Less Variable Costs Contribution margin What impact did the changes in number of detailings and average revenues (i.e., sales price) have on Classics, Ltd.s contribution margin? (Do not round intermediate calculations.) Contribution margi
Answer #1

Ans.

Actual (763 Units) Variable Cost Variance Sales Price Variance Flexible Budget (763 Units) Sales Activity Variance Master Budget (560 Units)
Revenue $71,200

$31,042 U

=($102,202 - $71,200)

$102,242= ($75,040/560 Units x 763 Units) $27,202 F =($102,242 - $75,040) $75,040
Less:
Variable Costs 35,520

$4,156 F = (39,676 - 35,520)

39,676= (29120/560units x 763 units) 10,556 U =(29,120 - 39,676) 29,120
Contribution Margin $35,680 $4,156 F $31,042 U $62,566 $16,646 F $45,920

Contribution Margin

Actual = $35,080

Budget = $45,920

Decreased by $10,240

Note : -

Actual (Based on actual activity of 763 unit detalings)

Flaxible budget (based on actual activity of 763 unit detalings)

Master budget (based on a production of 560 units detalings)

Last quarter Price = $75,040 / 560 = $134

$102,242 = Actual units x $134 = 763units X $134

Last quarter unit variable cost = 29,120 / 560 units = $52

$39,676 = Actual units X $52 = 763 x $52

The Numbar of detailings increased by 203 units which increased profit by $16,646 however the actual average price was ( $71,200 / 763 ) $93.32 so the average price per detailing by ($134 - $93.32 ) $40.68 As a result profit decreased by $10,240 ($45,920 - $35,680)